# Accurate stochastic simulation algorithm for multiscale models of infectious diseases

### v1

@@ 2.2.3 Implementation of accurate multiscale simulation
To simulate the within-host and population-level integrated model, we take the framework of a typical population-level Gillespie style SSA. The algorithm updates from event to event via a generalised next-reaction method (NRM) by storing all putative times for each unique event and implementing the event corresponding to the most imminent. Putative times for all Markovian events (in our case events ${(b)}$, ${(ds)}$, and ${(di)}$) are drawn in the same way as described in Subsection 2.1.2. However, times for the non-Markovian events (in our case ${(i)}$) need to be treated differently. We shall store in $\vec{\tau}$ a complete list of putative times for all Markovian events. In particular, $\vec{\tau}=\left(\tau^{(b)},\tau^{(ds)},\tau^{(di)}\right)^{\intercal}$. Furthermore, for each infected individual we also keep track of four pieces of information $\vec{\Psi}=\left\{\Psi_{j}\right\}_{j=1}^{i}$, $\vec{\psi}=\left\{\psi_{j}\right\}_{j=1}^{i}$, $\vec{\tau}^{(i)}=\left\{\tau_{j}^{(i)}\right\}_{j=1}^{i}$ and $\vec{\delta t}=\left\{\delta t_{j}\right\}_{j=1}^{i}$. Whilst it is the case that elements in $\vec{\Psi}$ are related to elements in $\vec{\delta t}$ in the same way as elements of $\vec{\Psi}+\vec{\psi}$ to those of $\vec{\delta t}+\vec{\tau}^{(i)}$, by the function $\Psi(\delta t)$, we find it convenient to store and track all four numbers for each infected individual.
It is necessary to begin by having a numerical solution for the within-host model and, together with (24), store $\Psi(\delta t)$ in a table. The table is stored with a discretisation of $\delta t$ given by $\delta$. Specifically, we store $\Psi(0)$, $\Psi(\delta)$, $\Psi(2\delta)$, and so on. Importantly, whilst this algorithm is accurate, the source of error is associated with the resolution of this table, so $\delta$ should be small. If $\delta$ is too small, however, using the look-up table becomes an expensive operation. From here on, we shall simply use the operators $\Psi$ and $\Psi^{-1}$ to reference using this look-up table, and, rounding to the closest stored value, if necessary. We use these operators liberally so that when they are applied to a vector of values, it is implied that the look-up table is used element by element, and therefore a vector of output values is being represented.
@@
We begin by initialising all population compartments ($s=s_{0}$, $i=i_{0}$). For each infected individual $j=1,\ldots,i_{0}$, it is necessary to determine an initial calendar time of infection $t_{j}$ and use this time to initialise $\vec{\delta t}$. We use the look-up table to initialise $\vec{\Psi}=\Psi(\vec{\delta t})$. Next we initialise all putative times. We initialise $\vec{\tau}$ by finding respective propensities for each element and sampling putative times using (11). We then initialise $\vec{\psi}$ for each element by using (27) and then use the look-up table to initialise $\vec{\tau}^{(i)}=\Psi^{-1}(\vec{\Psi}+\vec{\psi})-\vec{\delta t}$. The algorithm propagates forward by finding the most imminent event time $\tau^{(m)}=\mathrm{min}(\left\{\vec{\tau},\vec{\tau}^{(i)}\right\})$. After finding the event, time is updated and then the event $(m)$ is implemented (changing the compartments and deleting or initialising elements of $\vec{\Psi}$, $\vec{\psi}$, $\vec{\delta t}$ and $\vec{\tau}^{(i)}$, respectively) and then its putative time is re-sampled whilst all other affected putative times should be re-scaled (either using (13) if it is a Markovian event or (31) if it is a non-Markovian - infection - event). In particular, re-sampling and re-scaling of the non-Markovian infection events are described in Subsection 2.2.2 and involve updating respective elements of $\vec{\delta t}$, then $\vec{\Psi}$ followed by $\vec{\psi}$ and then $\vec{\tau}^{(i)}$. The algorithm then propagates until some pre-determined final time. As shorthand, we shall denote $\mathcal{Q}=\left\{\vec{\delta t},\vec{\tau}^{(i)},\vec{\Psi},\vec{\psi}\right\}$ as the collection of vectors that store information about each infected individual.
@@
2: Within-host model (for example (14)) to generate a tabulated function $\Psi(\delta t)$ with resolution $\delta$ as done in Equation (24);
@@
21:     else resample non-Markovian putative time for event $(m)$ using the method in Subsection 2.2.2.
@@
27:         update $\Psi_{j}$, rescale $\psi_{j}$, and then calculate the rescaled $\tau_{j}^{(i)}$ as described in Subsection 2.2.2.
@@ 3 Results
In this section, we aim to validate the accuracy of our novel algorithm (Algorithm 2) compared to an equivalent approximate time-driven method (Algorithm 1). In the approximate Algorithm 1, the calendar time is discretised with finite fixed time step $\Delta t$, while in our novel Algorithm 2, time discretisation is limited to the resolution of tabulated $\Psi(\delta t)$ associated with the within-host model. We shall denote this later discretisation as $\delta$ and note that it is possible to make this very small whilst incurring only the cost of reduced efficiency in using the table. It’s worth noting that whilst $\delta$ represents a coarse-graining of the within-host model dynamics, the calendar discretisation $\Delta t$ represents a course-graining in within-host and population scale models. That is, it is possible to miss intra-timestep population scale events with Algorithm 1 but not with Algorithm 2. However, if $\Delta t\rightarrow 0$, Algorithm 1 is expected to yield accurate results. Numerically, the $\Delta t_{\rm min}$ which guarantees the accuracy of the approximate algorithm is chosen such that
$\max_{t}\left\{\alpha^{(b)}+\alpha^{(ds)}+\alpha^{(di)}+s\sum_{j=1}^{i}v(t-t_{j})\right\}\Delta t\leq\epsilon:=0.01,$
@@
where $\max_{t}\left\{s\sum_{j=1}^{i}v(t-t_{j})\right\}\leq\max_{\delta t}\{v\}\times(s+i)^{2}/4$. For the initial condition $(S,I)=(800,80)$ people, we have $\Delta t_{\min}=9.37\times 10^{-5}$ day based on parameter values and within-host initial conditions shown in Tables 12. To assess the accuracy of our novel algorithm, we contrast consistent discretisation $\delta=\Delta t$ between the two algorithms and vary this discretisation geometrically over 131 points. Specifically, $\delta=\Delta t=\Delta t_{\min}\times 1.084^{(0:130)}$ days. In Figure 5, we visualise the course-graining of the within-host viral load $v$ as a result of using $\delta_{\min}=9.37\times 10^{-5}$ day, $\delta=0.28$ day, and $\delta=1$ day under the assumption that within-host virus dynamics occur as a result of specific model parameters presented in Table 1. We include in this figure the discretisation for both Algorithm 1 and 2. The reason for this is subtle. Sampled viral loads under Algorithm 2 are tabulated starting at $\delta t=0$ day. On the other hand, we assume that in Algorithm 1 that an infection begins on a time step and thus we sample halfway between time steps as the best approximation of the viral load on this time interval. The difference in implied numerical viral load when $\Delta t=\delta=1$ day is clear in Figure 5. We refer to the error associated with the discretisation of the within-host model as ‘resolution error’. We are not deterred from labelling Algorithm 2 as being an ‘exact’ algorithm since the algorithm is, at least in principle, exact if the continuous function $\Psi(\delta t)$ is known in the same sense that the Gillespie algorithm is exact despite the fact that there may be numerical error in the propensities.
@@
Figure 5: The resolution of harvesting the within-host viral load with $\Delta t=\delta$. The within-host system, introduced in (14) in Subsection 2.1.3, is solved using MATLAB’s $ode15s$ solver with both absolute and relative tolerances set to $10^{-8}$. The initial conditions and the parameter values for the within-host system are listed in Table 1.
@@ 4 Discussion
When simulating a population-scale model of the spread of a disease, it is not uncommon to see the implementation of autonomy (memorylessness) in the model. This is the case with ODE and also Gillespie-type stochastic simulations. In the case of small populations, where the outcome of an epidemic may hinge on the infections of one or two infected individuals, these memoryless models do not factor in the progress the individuals are making with the disease. To track an individual’s viral load, a within-host model is required to be coupled to the population-scale model. We have shown that the introduction of a within-host model for each individual can be achieved with extremely small time steps. However, such a time-driven algorithm becomes inaccurate (Figures 67) if the fixed time step is no longer small enough. It is important that the time step chosen is much smaller than the natural rate of change of viral load for an infected individual. Additionally, since working with predefined finite time steps, it is important to keep the time step much smaller than the inverse of the rate of infection and therefore the time step should depend on the unknown future population-level and the overall within-host level dynamics, which is almost impossible to estimate. This is because it is assumed that the probability for an event to occur in any given time step should be small. Whilst Tau-leaping style approaches might also be useful, they too require some consideration for the rate at which events are occurring in order to retain sufficiently high accuracy. In this paper, we have sidestepped this challenge and developed a novel stochastic algorithm (Algorithm 2) for multiscale systems which incorporates a within-host viral load dynamic into the infection rates of individuals embedded in a population-scale model of an infectious disease. The key strength of our algorithm lies in its accuracy and generality. Specifically, event times are simulated in a similar way to the Gillespie algorithm but taking into consideration the explicit time dependence in individual viral loads. We conclude from Section 3 that as long as the within-host data are extracted at a sensible resolution, the simulated results will always be accurate using our approach. Additionally, while we have introduced the concepts of this novel approach using simple within-host and population-level models, our algorithm can be applied to similar compartmental models at both levels. We thus conclude that our accurate Algorithm 2 can be broadly applied across the field of multiscale infectious disease modelling.
@@
The primary contribution of this study lies in the development of a novel algorithm and the demonstration of its accuracy. The evaluation of its efficiency is reserved for future investigation. This novel algorithm can also be extended in several directions. For instance, we can explore scenarios where the whole population still shares a single within-host system, but it is no longer deterministic and includes noise. Alternatively, given a single deterministic within-host system, we can investigate situations where different infectious individuals have varying within-host time scales. Moreover, our algorithm can be expanded to accommodate heterogeneous populations, such as those with age structures, where each sub-population shares a single deterministic within-host system. In this case, one can simply create different lookup tables (see (24)) that tabulate different within-host dynamics for each sub-population.
@@
Our work also holds broader implications and significance in real-world scenarios. Until now, while considerable effort has been dedicated to nesting the within-host system into the population-level system in multiscale modelling of infectious diseases [TAKT23, MTSM15, MAD08], the reverse direction has not been fully explored [FVHTSL12]. The novel accurate Algorithm 2 we have developed can serve as an ideal tool for generating and harvesting global within-host information averaged across the entire population. In this sense, one may uncover potential relationships between these two subsystems and, consequently, nest the population-level system back into the within-host one. The advantage of doing so lies in the fact that, after coarse-graining, a single set of ODEs can be achieved, incorporating both population-level and averaged global within-host dynamics. As a result, it would be much more efficient to test, refine, and calibrate the simple yet comprehensive coarse-grained ODE model to real-life data. Parameter inference would also be more realistic, enabling predictions and the formulation of policies to slow down disease transmission.
5 Data accessibility

### v2

@@ 2.2.3 Implementation of accurate multiscale simulation
To simulate the within-host and population-level integrated model, we take the framework of a typical population-level Gillespie-style NRM. The algorithm updates from event to event by storing all putative times for each unique event and implementing the event corresponding to the most imminent. Putative times for all Markovian events (in our case events ${(b)}$, ${(ds)}$, and ${(di)}$) are drawn in the same way as described in Section 2.1.2. However, times for the non-Markovian events (in our case ${(i)}$) need to be treated differently. We shall store in $\vec{\tau}$ a complete list of putative times for all Markovian events. In particular, $\vec{\tau}=\left(\tau^{(b)},\tau^{(ds)},\tau^{(di)}\right)^{\intercal}$. Furthermore, for each infected individual we also keep track of four pieces of information $\vec{\Psi}=\left\{\Psi_{j}\right\}_{j=1}^{i}$, $\vec{\psi}=\left\{\psi_{j}\right\}_{j=1}^{i}$, $\vec{\tau}^{(i)}=\left\{\tau_{j}^{(i)}\right\}_{j=1}^{i}$ and $\vec{\delta t}=\left\{\delta t_{j}\right\}_{j=1}^{i}$. Whilst it is the case that elements in $\vec{\Psi}$ are related to elements in $\vec{\delta t}$ in the same way as elements of $\vec{\Psi}+\vec{\psi}$ to those of $\vec{\delta t}+\vec{\tau}^{(i)}$, by the function $\Psi(\delta t)$, we find it convenient to store and track all four numbers for each infected individual.
To implement our method, we first obtain a numerical solution for the within-host model and, following (23), store $\Psi$ in a table. For efficiency, as we explain later, we instead construct two look-up tables. The first table, $\Psi(\delta t)$, is stored using a uniform discretisation of $\delta t$ with step size $\delta$, while the second table, $\delta t(\Psi)$, has the same length as $\Psi(\delta t)$ but is discretised with respect to $\Psi$ using step size $\delta\Psi$. Importantly, whilst this approach is accurate, the primary source of error stems from the resolution of these tables, necessitating sufficiently small $\delta$ and $\delta\Psi$. Then, to evaluate $\Psi$ for an infectious individual $j$ with age of infection $\delta t_{j}$, we retrieve its value from $\Psi(\delta t)$ by indexing the table at $\delta t_{j}/\delta$, rounded to the nearest integer. Similarly, evaluating $\Psi^{-1}$ follows the same principle by accessing $\delta t(\Psi)$ at $\psi_{j}/\delta\Psi$, also rounded to the nearest integer. From now on, we denote these look-up operations as $\Psi$ and $\Psi^{-1}$, respectively. When applied to a vector, these operations are performed element-wise, yielding a corresponding vector of output values. A key advantage of this approach is its computational efficiency. Both $\Psi$ and $\Psi^{-1}$ operations involve only floating-point division and array indexing, which are computationally inexpensive, regardless of the resolution of $\delta$ and $\delta\Psi$ or the size of the tables $\Psi(\delta t)$ and $\delta t(\Psi)$. This highlights a crucial benefit of our novel exact algorithm – high efficiency is achieved alongside accuracy.
@@
We begin by initialising all population compartments ($s=s_{0}$, $i=i_{0}$). For each infected individual $j=1,\ldots,i_{0}$, it is necessary to determine an initial calendar time of infection $t_{j}$ and use this time to initialise $\vec{\delta t}$. We use the look-up tables to initialise $\vec{\Psi}=\Psi(\vec{\delta t})$. Next we initialise all putative times. We initialise $\vec{\tau}$ for Markovian events ($(b)$, $(ds)$, $(di)$) by finding respective propensities for each element and sampling putative times using (11). We then initialise $\vec{\psi}$ for each element by using (26) and then use the look-up tables to initialise $\vec{\tau}^{(i)}=\Psi^{-1}(\vec{\Psi}+\vec{\psi})-\vec{\delta t}$ for the non-Markovian event $(i)$. The algorithm propagates forward by finding the most imminent event time $\tau^{(m)}=\mathrm{min}(\left\{\vec{\tau},\vec{\tau}^{(i)}\right\})$. After finding the event, time is updated and then the event $(m)$ is implemented (changing the compartments and deleting or initialising elements of $\vec{\Psi}$, $\vec{\psi}$, $\vec{\delta t}$ and $\vec{\tau}^{(i)}$, respectively) and then its putative time is re-sampled whilst all other affected putative times should be re-scaled (either using (12) if it is a Markovian event or (30) if it is a non-Markovian – infection – event). In particular, re-sampling and re-scaling of the non-Markovian infection events are described in Section 2.2.2 and involve updating respective elements of $\vec{\delta t}$, then $\vec{\Psi}$ followed by $\vec{\psi}$ and then $\vec{\tau}^{(i)}$. The algorithm then propagates until some pre-determined final time. As shorthand, we shall denote $\mathcal{Q}=\left\{\vec{\delta t},\vec{\tau}^{(i)},\vec{\Psi},\vec{\psi}\right\}$ as the collection of vectors that store information about each infected individual.
The pseudocode for this algorithm is shown in Algorithm 2, where the MATLAB implementation and all data generated from the simulations are available on Github at https://github.com/YuanYIN99/Accurate-stochastic-simulation-algorithm-for-multiscale-models-of-infectious-diseases.git.
@@
2: Within-host model (for example (13)) to generate two look-up tables $\Psi(\delta t)$ and $\delta t(\Psi)$ as described in Section 2.2.3;
@@
21:     elseresample non-Markovian putative time for event $(m)$ using the method in Section 2.2.2.
@@
27:         update $\Psi_{j}$, rescale $\psi_{j}$, and then calculate the rescaled $\tau_{j}^{(i)}$ as described in Section 2.2.2.
@@ 3 Results
We now evaluate the performance of our novel algorithm by comparing its accuracy and efficiency against the time-driven approximate method. We focus on a test case defined by the within-host dynamics in Table 1 and the population-level parameters in Table 2. In this section, we investigate how varying the initial population sizes ($(S,I)=(100,10)$, $(200,20)$, $(400,40)$, and $(800,80)$ people) and the frequencies at which within-host information is harvested affects the performance of both algorithms. Each simulation is performed with $800$ repetitions using parallel computing, and all simulations are executed on a local computer with a 14-core CPU.
@@
3.1 Accuracy
Figure 5: The resolution of harvesting the within-host viral load using a constant time step $\Delta t=\delta$ and a constant probability step $\delta\Psi$. Note that $\delta\Psi$ is defined so that the within-host system is extracted at the same resolution as $\delta$. The within-host system, introduced in (13) in Section 2.1.3, is solved using MATLAB’s $ode15s$ solver with both absolute and relative tolerances set to $10^{-8}$. The initial conditions and the parameter values for the within-host system are listed in Table 1.
@@ 3.1 Accuracy
In this section, we aim to validate the accuracy of our novel algorithm (Algorithm 2) compared to the approximate time-driven method (Algorithm 1). In the approximate Algorithm 1, the calendar time is discretised with finite fixed time step $\Delta t$, while in our novel Algorithm 2, time discretisation is limited to the resolution of tabulated $\Psi(\delta t)$ thus $\delta t(\Psi)$ associated with the within-host model. Recall that since the length of the latter look-up table $\delta t(\Psi)$ is equal to that of the former $\Psi(\delta t)$, the discretisation of $\Psi$ in the inverse table is determined by the time discretisation $\delta$. We therefore characterise the discretisation of both look-up tables in the novel Algorithm 2 as $\delta$. Note that it is possible to make $\delta$ very small without incurring the cost of reduced efficiency in using these tables (see Section 2.2.3). It’s worth noting that whilst $\delta$ represents a coarse-graining of the within-host model dynamics, the calendar time discretisation $\Delta t$ represents a course-graining in within-host and population scale models. That is, it is possible to miss intra-timestep population scale events with Algorithm 1 but not with Algorithm 2. However, if $\Delta t\rightarrow 0$, Algorithm 1 is expected to yield accurate results. Numerically, the $\Delta t_{\rm min}$ which guarantees the accuracy of the approximate algorithm is chosen such that
$\max_{\Delta t}\left\{\alpha^{(b)}+\alpha^{(ds)}+\alpha^{(di)}+s\sum_{j=1}^{i}v(t-t_{j})\right\}\Delta t\leq\epsilon:=0.01,$
where $\max_{\Delta t}\left\{s\sum_{j=1}^{i}v(t-t_{j})\right\}\leq\max_{\Delta t}\{v\}\times(s+i)^{2}/4$. For the initial condition $(S,I)=(800,80)$ people, we have $\Delta t_{\min}=9.37\times 10^{-5}$ day based on parameter values and within-host initial conditions shown in Tables 1-2. To assess the accuracy of our novel algorithm, we contrast consistent discretisation $\delta=\Delta t$ between the two algorithms and vary this discretisation geometrically over $131$ points. Specifically, $\delta=\Delta t=\Delta t_{\min}\times 1.084^{(0:130)}$ days.
@@
In Figure 5, we visualise the coarse-graining of the within-host viral load $V$ using constant time steps $\delta=9.37\times 10^{-5}$ days, $\delta=0.28$ days, and $\delta=1$ day, along with the corresponding constant probability steps, assuming that the within-host virus dynamics follow the model parameters specified in Table 1. Recall that constant time tabulation is used in the time-driven Algorithm 1 as well as in the computation of $\Psi(\delta t)$ in our novel exact Algorithm 2, whereas constant probability tabulation is employed for computing $\Psi^{-1}$ in Algorithm 2. We assume that an infection begins on a time step and thus we sample halfway between time steps as an approximation of the viral load on this time interval. We refer to the error associated with the discretisation of the within-host model as ‘resolution error’. We are not deterred from referring to Algorithm 2 as an exact algorithm, since it is, at least in principle, exact – provided that the continuous function $\Psi(\delta t)$ (and consequently $\Psi^{-1}$) is known – just as the Gillespie algorithm is considered exact despite potential numerical errors in the evaluation of propensities.
@@
3.2 Efficiency
Figure 8: Average runtime (excluding within-host tabulation time) of the approximate time-driven Algorithm 1 and the novel Algorithm 2 for simulations shown in Figure 7. The orange, green, blue, and light purple curves represent different initial conditions: $(S,I)=(100,10)$, $(200,20)$, $(400,40)$, and $(800,80)$ people, respectively. The markers indicate the maximum value of $\delta=\Delta t$ that generates accurate results, with a relative error bounded by $0.05$. The initial conditions and parameter values for the within-host system are provided in Table 1, while the population-level model parameter values are listed in Table 2.
In this section, we aim to validate the efficiency of our novel algorithm (Algorithm 2) compared to the time-driven method (Algorithm 1). The primary computational cost in the approximate time-driven algorithm arises from tabulating the force of infection $v$ at the start of the simulation, which has a complexity of $O(1/\Delta t)$, and determining the next event. For the latter, given a calendar time discretisation of $\Delta t$, we must sum over $I$ forces of infection at each time step, leading to a time complexity of $O(I/\Delta t)$. Consequently, the overall complexity of the approximate time-driven algorithm is $O(I/\Delta t)$. In contrast, the computational complexity of the novel exact algorithm primarily stems from two factors: (1) constructing two look-up tables of resolution $\delta$ at the start of the simulation and (2) determining event-driven times while updating the within-host systems of infectious individuals. Construction of the two look-up tables has a complexity of $O(1/\delta)$, while the latter operation is $O(I)$, as it only involves floating-point divisions and array indexing for each infectious individual. If a total of $M$ events occur before the simulation end time, $t_{\text{end}}$, the overall time complexity of our exact algorithm is $O(MI+1/\delta)$. An important observation is that, due to the rejection nature of the time-driven algorithm, the number of events satisfies $M\leq t_{\text{end}}/\Delta t$. Therefore, with resolutions $\Delta t=\delta$, the time-driven algorithm exhibits no better time complexity compared to our novel exact algorithm, when the tabulation time is excluded (as it’s a fixed cost).
We then compare the efficiency of the novel exact Algorithm 2 with the approximate time-driven Algorithm 1, across varying population sizes and resolutions for extracting within-host information. Figure 8 shows the average runtime per repetition, excluding the time required to tabulate the within-host system. We observe that the runtime of Algorithm 1 increases with finer time steps ($\Delta t$) and bigger population sizes. This observation is consistent with the previous theoretical complexity analysis. On the other hand, the runtime of Algorithm 2 is independent of the resolution used to tabulate within-host information, but increases with population size, reflecting the greater number of events occurring over the simulation time frame. Additionally, the runtime of Algorithm 2 appears less smooth due to stochastic variability in the number of events per simulation. We now focus on regions where the time discretisations ($\Delta t$ and $\delta$) are smaller than those highlighted by the markers in Figure 8, as these indicate the critical resolutions required to ensure accuracy (see Section 3.1). In practice, we find that the novel exact algorithm is more efficient – or at least equivalent – to the time-driven approach in these regimes. We also note that, in this specific test case, the tabulation time dominates the total runtime. However, this need not be the case in other scenarios – for example, when the within-host system stabilises rapidly.
@@ 4 Discussion
When simulating a population-scale model of the spread of a disease, it is not uncommon to see the implementation of autonomy (memorylessness) in the model. This is the case with ODE and also Gillespie-type stochastic simulations. In small populations, where the outcome of an epidemic may hinge on the infections of one or two infected individuals, these memoryless models fail to account for the stages of their infection. To track an individual’s viral load, a within-host model is required to be coupled to the population-scale model. We have shown that the introduction of a within-host model for each individual can be achieved with extremely small time steps. However, such a time-driven algorithm becomes inaccurate (Figures 6-7) if the fixed time step is no longer small enough. It is important that the time step chosen is much smaller than the natural rate of change of viral load for an infected individual. Additionally, since working with predefined finite time steps, it is important to keep the time step much smaller than the inverse of the rate of infection and therefore the time step should depend on the unknown future population-level and the overall within-host level dynamics, which is almost impossible to estimate. This is because it is assumed that the probability for an event to occur in any given time step should be small. Whilst Tau-leaping style approaches might also be useful, they too require some consideration for the rate at which events are occurring in order to retain sufficiently high accuracy.
In this paper, we have sidestepped this challenge and developed a novel stochastic algorithm (Algorithm 2) for multiscale systems which incorporates a within-host viral load dynamic into the infection rates of individuals embedded in a population-scale model of an infectious disease. The key strength of our algorithm lies in its accuracy, efficiency, and generality. Specifically, event times are simulated in a similar way to the Gillespie algorithm but taking into consideration the explicit time dependence in individual viral loads. We conclude from Section 3 that as long as the within-host data are extracted at a sensible resolution, the simulated results will always be accurate using our approach. Furthermore, extracting within-host information at finer resolutions does not compromise efficiency (Figure 8). While we have introduced the concepts of this novel approach using simple target cell-limited within-host and $SI$ population-level models, our algorithm can be applied to similar compartmental models at both levels. For example, for a different within-host model, the primary change would be in the look-up table $\Psi$ ((23) in Section 2.2), which records the within-host quantities that influence the population-level dynamics. For a different population-level model, we would need to consider different events, such as infection, recovery, and waning immunity. However, the method for sampling non-Markovian event putative times using the look-up table, resampling and rescaling those times after an event, remains unchanged. Therefore, we conclude that our accurate Algorithm 2 can be broadly applied across the field of multiscale infectious disease modelling.
@@
This novel algorithm can also be extended in several directions. For instance, we can explore scenarios where the whole population still shares a single within-host system, but it is no longer deterministic and includes noise. Alternatively, we can build a heterogeneous population. For instance, in age-structured populations where each sub-population shares a single deterministic within-host system, one can simply create different look-up tables that tabulate different within-host dynamics for each sub-population. Moreover, we can extend our algorithm to investigate situations where different infectious individuals have varying within-host time scales and amplitudes by tabulating a dimensionless version of the solution of the within-host model. However, we emphasise that our algorithm cannot be extended to scenarios where each individual has a unique set of within-host system. In such cases, it would be extremely challenging for any simulation approach to generalise.
@@
Our work also holds broader implications and significance in real-world scenarios. Until now, considerable effort has been dedicated to nesting the within-host system into the population-level one in multiscale modelling of infectious diseases [TAKT23, MTSM15, MAD08]. However, the reverse direction – coupling population-level dynamics into the within-host system – has not been fully explored [FVHTSL12]. A key step in investigating this reverse nesting is obtaining a global within-host system by averaging over the entire population. The novel and accurate Algorithm 2 we have developed is an ideal tool for generating and extracting such information. This allows for uncovering potential relationships between population-level dynamics and global within-host parameters, enabling the derivation of a single set of ODEs that incorporates dynamics at both the population and global within-host levels after coarse-graining. As a result, it would be much more efficient to test, refine, and calibrate the simple yet comprehensive coarse-grained ODE model to real-life data. Parameter inference would also be more realistic, enabling predictions and the formulation of policies to slow down disease transmission.
5 Code and data accessibility